OpenStudy (andrewthecookie):

50 is 6.25% of what number? One last problem for @AriPotta

OpenStudy (andrewthecookie):

Am I right

OpenStudy (aripotta):

i believe you are correct :)

TheSmartOne (thesmartone):

He is correct...

OpenStudy (andrewthecookie):

And one last problem because on the one 6% of $179,000 dollars I misread and it says write a proportion and show my work. So what would be the proportion

OpenStudy (andrewthecookie):

@TheSmartOne Can answer?

TheSmartOne (thesmartone):

wouldn't it be \(\Huge \frac{6}{100}=\frac{x}{179,000}\)

OpenStudy (aripotta):

\[\frac{ part }{ whole }\frac{ 6 }{ 100 }=\frac{ x }{ 179000 }\]

TheSmartOne (thesmartone):

^ so yeah :D

OpenStudy (andrewthecookie):

Ok thx guys!!

TheSmartOne (thesmartone):

No problem :P

OpenStudy (andrewthecookie):

:D Hopefully i get all of the questions

OpenStudy (aripotta):

no problemo :)

OpenStudy (andrewthecookie):

I wish I could medal two people

TheSmartOne (thesmartone):

Your medals got me to 97!!!

OpenStudy (aripotta):

yey

OpenStudy (andrewthecookie):

:o Cool :D

OpenStudy (aripotta):

lol thesmartone, just a few days ago, you were at like 92. you're zooming by

TheSmartOne (thesmartone):

I know :P

TheSmartOne (thesmartone):

99SS here I come

OpenStudy (andrewthecookie):

Satellite has 100 SS

TheSmartOne (thesmartone):

Only 1 person can have it...

OpenStudy (andrewthecookie):

So he's special >:O

OpenStudy (aripotta):

he's beyond special

TheSmartOne (thesmartone):

^^ lol